OpenStudy (anonymous):

~~~~~GIVING OUT A MEDAL AND FANNING ANYONE WHO HELPS ME THANK YOU ^_^!!!!~~~~~ Factor completely: 2x3 + 10x2 + 4x + 20 2(x3 + 5x2 + 2x + 10) 2[(x + 5)(x2 + 2)] (2x + 10)(x2 + 2) (x + 5)(2x2 + 4)

OpenStudy (anonymous):

@jazzyfa30

OpenStudy (nory):

First, there is a common factor among all terms. Do you see why we can factor out a 2 from everything?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

The two can go into any of those

OpenStudy (anonymous):

I think

OpenStudy (nory):

2(x^3 + 5x^2 + 2x + 5)

OpenStudy (anonymous):

let me check if it's right

OpenStudy (uri):

2x3 + 10x2+ 4x + 20 What's common in this? 2(x^3+5x^2+2x+10 )

OpenStudy (anonymous):

I'm going to make an educated guess because i'm just learning this but I believe you divided everything by two

OpenStudy (uri):

Ya..

OpenStudy (anonymous):

So that's the answer?

OpenStudy (saifoo.khan):

Unfortunately not.

OpenStudy (anonymous):

Oh... where did I go wrong?

OpenStudy (anonymous):

-5,1.41,-1.43

OpenStudy (anonymous):

? whats that?

OpenStudy (anonymous):

-5,1.41,-1.43

OpenStudy (anonymous):

@Rezwanahmed and @50_cent u have solved the cubic equation which is not mentioned here

OpenStudy (anonymous):

value of x

OpenStudy (uri):

I just solved it.It's A.

OpenStudy (anonymous):

Lol @saifoo.khan said it was wrong but I need an explanation for how its wrong

OpenStudy (saifoo.khan):

That wasn't completely factorized.

OpenStudy (anonymous):

@HelloGoodmorning as there is no common coming out.use alternative method.

OpenStudy (jazzyfa30):

the 2nd one

OpenStudy (jazzyfa30):

2[(x + 5)(x2 + 2)]

OpenStudy (anonymous):

Oh ok @saifoo.khan and @heybhai what is the alternative method and @jazzyfa30 can you explain how you got the answer to that

OpenStudy (anonymous):

Oh ok

OpenStudy (uri):

2(x3 + 5x2 + 2x + 10) 2[x^2(x+5) +2(x+5) 2(x+5)(x^2+2)

OpenStudy (anonymous):

(2x+10)(x^2+2)

OpenStudy (anonymous):

so @Rezwanahmed thats the answer?

OpenStudy (uri):

B is the answer.

OpenStudy (saifoo.khan):

Answer is E.

OpenStudy (jazzyfa30):

no it isn't it is B

OpenStudy (anonymous):

e? @saifoo.khan

OpenStudy (anonymous):

value of x

OpenStudy (anonymous):

@Rezwanahmed what value of x?

OpenStudy (anonymous):

@HelloGoodmorning if u wanna solve this assuming this as an equation then solution for any cubic equation ax^3+bx^2+cx+d=0. assume M,N and O as M+N+O= -b/a MN+NO+OM=c/a MNO=-d/a solve them and get the solution

OpenStudy (jazzyfa30):

dude u must be crazy think im boutta sit here and solve that i got my own math work to do like i said the answer is B and that's that

OpenStudy (anonymous):

(2X+10)(x2+2)I think it is answer

OpenStudy (anonymous):

I think so too @Rezwanahmed

OpenStudy (anonymous):

Its a pretest im not to worried i dont know why they did a pretest before you learn the lesson but it still counts as a grade